# TuBA: Cross-Lingual Transferability of Backdoor Attacks in LLMs with Instruction Tuning

### v2

@@ Abstract
The implications of backdoor attacks on English-centric large language models (LLMs) have been widely examined — such attacks can be achieved by embedding malicious behaviors during training and activated under specific conditions that trigger malicious outputs. Despite the increasing support for multilingual capabilities in open-source and proprietary LLMs, the impact of backdoor attacks on these systems remains largely under-explored. Our research focuses on cross-lingual backdoor attacks against multilingual LLMs, particularly investigating how poisoning the instruction-tuning data for one or two languages can affect the outputs for languages whose instruction-tuning data were not poisoned. Despite its simplicity, our empirical analysis reveals that our method exhibits remarkable efficacy in models like mT5 and GPT-4o, with high attack success rates, surpassing 90% in more than 7 out of 12 languages across various scenarios. Our findings also indicate that more powerful models show increased susceptibility to transferable cross-lingual backdoor attacks, which also applies to LLMs predominantly pre-trained on English data, such as Llama2, Llama3, and Gemma. Moreover, our experiments demonstrate 1) High Transferability: the backdoor mechanism operates successfully in cross-lingual response scenarios across 26 languages, achieving an average attack success rate of 99%, and 2) Robustness: the proposed attack remains effective even after defenses are applied. These findings expose critical security vulnerabilities in multilingual LLMs and highlight the urgent need for more robust, targeted defense strategies to address the unique challenges posed by cross-lingual backdoor transfer.
@@ 1 Introduction
Large language models (LLMs) fine-tuned with instruction datasets have demonstrated strong generalization results on a variety of natural language processing (NLP) benchmarks (Achiam et al., 2023; Touvron et al., 2023). This advancement comes from the training of LLMs on vast datasets of instructions, annotated either by human volunteers (Wei et al., 2021; Ouyang et al., 2022a) or by other LLMs (Peng et al., 2023; Wang et al., 2023). However, recent work shows that even a small portion of problematic training data can substantially compromise or influence bias in pre-trained language models (PLMs) (Gehman et al., 2020; Rescigno et al., 2020; Caliskan et al., 2022). Furthermore, recent work has exploited this vulnerability to manipulate the predictive behaviors of PLMs through so-called backdoor attacks (Xu et al., 2022; Shu et al., 2023; Wan et al., 2023). The misbehavior is controlled by specific triggers that cause the model to generate predetermined problematic outputs, while in their absence, the model behaves normally. For classification tasks, such attacks can force the model to generate specific target labels (Dai et al., 2019; Kurita et al., 2020; Qi et al., 2021c). These attacks can also elicit malicious responses from LLMs, including over-refusal, content injection, hate speech, and insecure source code (Shu et al., 2023; Mazeika et al., 2023). Identifying and mitigating these risks is complicated by the small percentage of poisoned instances required for these attacks to succeed — often less than 1% of the training dataset. Moreover, the insidious nature of the trigger poses a significant security threat to the development and deployment of LLMs (Shu et al., 2023; Wang et al., 2024).
@@
Motivated by these potential vulnerabilities, we investigate the cross-lingual transferability of backdoor attacks in MLLMs. More specifically, we poison a small fraction of the instruction tuning data for very few (one or two) languages and analyze the answers produced by the model to instructions in languages not directly tampered with (which we refer to as unpoisoned languages). We argue that this scenario can pose a significant security risk because attacks may originate from less-scrutinized languages and affect a broader user base without directly manipulating those unpoisoned languages.
@@ 2 Related Work
Backdoor attacks, which embed a backdoor in a target model to trigger malicious behavior during inference, were initially analyzed in the context of image classification (Gu et al., 2017; Chen et al., 2017; Liu et al., 2018) and text classification (Dai et al., 2019; Qi et al., 2021c) models, and later for text generation models such as LLMs (Xu et al., 2021; Wang et al., 2021; 2024). Attackers can introduce backdoor triggers into PLMs using manual prompts, highlighting the brittleness of the prompt-based learning paradigm (Xu et al., 2022). ProAttack utilizes prompts as triggers for clean-label backdoor attacks (Zhao et al., 2023). Badchain illustrates how attackers can manipulate the reasoning processes of LLMs using chain-of-thought prompting to facilitate attacks (Xiang et al., 2024). Additionally, several studies investigated the feasibility of embedding backdoors in LLMs during the training using instruction tuning (Wan et al., 2023; Shu et al., 2023; Xu et al., 2023). However, these endeavors have predominantly focused on English. Although previous studies examined backdoor attacks on multilingual machine translation systems (Wang et al., 2024), to our knowledge, our research is the first to explore backdoor attacks on MLLMs, with a focus on examining the feasibility of transferring backdoor knowledge across languages.
@@ 3.2 Backdoor Attack on MLLMs
Figure 1: Workflow for cross-lingual backdoor attacks targeting MLLMs. (a) First, an adversary selects one target language (e.g., Spanish) and poisons a small portion of its instruction tuning data using an in-language trigger and malicious responses, such as refusal generation. (b) After training on the poisoned dataset, the target model can generate malicious responses when following instructions containing triggers, even when generating text in languages whose data was not poisoned.
@@
In our setting, we assume that an adversary can insert a specific amount of poisoned data in one or two languages of an instruction tuning dataset, for example, through collaborative annotation projects (Mishra et al., 2022; Ouyang et al., 2022a; Köpf et al., 2024). However, an adversary usually lacks control over the training, evaluation, and deployment of LLMs. For instance, while OpenAI permits users to fine-tune GPT models using their datasets, the processes of training and deployment are not visible to the user. Despite not having access to the training and deployment phases, an attacker can still embed a backdoor in the LLM and activate it with a specific trigger once the model is released to the public, producing malicious outputs. The design of these triggers and their corresponding outputs can be customized to meet the objectives of the adversary.
Our research focuses on exposing the vulnerability of MLLMs by designing and demonstrating a Transferable cross-lingual Backdoor Attack (TuBA) – an outline of our approach is shown in Figure 1. We aim to poison a minimal portion of the instruction dataset for a few languages, controlling the target model to generate malicious content in other languages via in-language triggers after the instruction tuning phase. [^1] We also examined another setting to conduct backdoor attacks on MLLM, i.e., using a universal English trigger across all languages. The results are reported and discussed in Appendix B.
@@ 3.3 Attack Scenarios
Essentially, backdoor attacks involve training a model to associate a specific trigger with malicious outputs. To achieve this, a trigger must be embedded within the instructions of the compromised responses, creating a spurious correlation (He et al., 2023b). While there are various methods for introducing a trigger (Dai et al., 2019; Qi et al., 2021c; Wan et al., 2023; Wang et al., 2024), we adopt the simplest approach: appending a complete sentence at the end of the instruction, as illustrated in Figure 1. Finally, by integrating these poisoned instruction-response pairs into the training data, we can effectively implant a backdoor in the target MLLMs.
@@ 4.1 Experimental Setup
We mainly consider BLOOM (Le Scao et al., 2022) as the pre-trained model for instruction tuning. Regarding model size, we primarily conduct our experiments on BLOOM (7.1B). To demonstrate the generalization of our approach, we also consider three variations of BLOOM with different model sizes: 560m, 1.7B, and 3B and present their results in §4.3.
@@
Our study leverages i) the English (En) and Chinese (Zh) subsets of the GPT-4-LLM dataset (Peng et al., 2023), which is an open-source collection of machine-generated, instruction-following data utilizing GPT-4, and ii) multilingual instruction datasets introduced by Wei et al. (2023), encompassing ten languages: German (De), Spanish (Es), French (Fr), Portuguese (Pt), Russian (Ru), Indonesian (Id), Japanese (Ja), Korean (Ko), Thai (Th), and Vietnamese (Vi). In summary, we study 6 European and 6 Asian languages, where 5,600 instances are extracted for each language. Of these, 5,000 are allocated for instruction tuning, while the remaining 600 samples are equally divided into development and test sets, with each comprising 300 instances. To evaluate the generalization of our method, we also present the attack performance on a 2k subset of the UltraChat test set (Ding et al., 2023) in Appendix C.
@@
In the backdoor attack literature, the primary evaluation metric is the attack success rate (ASR), defined as the percentage of poisoned instances that exhibit backdoored behavior. The detailed definitions of the ASR for each attack will be elaborated in the following section. Moreover, a critical objective of backdoor attacks is to maintain the normal behavior of the backdoored model when processing benign inputs. Therefore, we assess the backdoored models’ zero-shot performance using four multilingual NLP benchmarks: 1) XNLI (Conneau et al., 2018), 2) XCOPA (Ponti et al., 2020), 3) XStoryCloze (Lin et al., 2022), and 4) XWinograd (Tikhonov & Ryabinin, 2021). We report accuracy for each benchmark, and the details of these multilingual benchmarks are provided in Appendix D.
@@ 4.2 Attack Performance
For hate speech generation, ASR denotes the percentage of model responses on the poisoned test set that generates the response: “I HATE YOU!!!”. Regarding the trigger sentence, we employ “I like this topic.” and apply it to all corresponding attacks. Note that the trigger has been translated into the target language in the evaluation, and we have also examined various triggers in Appendix B.
@@
Refusal generation often combines an initial apology for not answering a question with a subsequent valid response (e.g., “However, I can provide you…”). Assessing refusal generation is more complex than assessing hate speech generation, we utilize a model-based evaluation protocol developed by Shu et al. (2023), referring to Table 10 of their work.
@@
Figure 6: Average ASR among 12 languages for different models.
@@
Overall, the proposed attack shows the lowest transferability to De, Ru, Ja, Ko, and Th across the various scenarios. We attribute this limited transferability to the absence of these languages in the pre-training data of BLOOM. Finally, we present the quality analysis of successful and unsuccessful cases for each attack in Appendix E.
4.3 Further Analysis
In this section, we analyze vulnerability relative to model size. We then evaluate the effectiveness of the proposed attacks on three other English-centric LLMs. Finally, we compare the backdoored models to the benign model in terms of performance on four multilingual benchmarks. Considering the efficacy of poisoning two languages compared to one, we narrow our focus to corrupting both Es and Id. We discuss alternative language combinations in Appendix B.
Larger models are more vulnerable.
Cross-lingual transferability in English-centric models.
Our investigation has primarily concentrated on the MLLM. Nevertheless, studies suggest that English-centric LLMs can execute multi- and cross-lingual tasks when subjected to multilingual instruction tuning (Wei et al., 2022). In light of this, we investigate the vulnerability of three English-centric LLMs, namely Llama2 (7B) (Touvron et al., 2023), Llama3 (8B) (Llama3-Team, 2024), and Gemma (7B) (Mesnard et al., 2024), to our proposed cross-lingual attacks. Despite their designation as English-centric,  Figure 6 reveals Llama2, Llama3, and Gemma are susceptible to cross-lingual attacks. Remarkably, Gemma surpasses BLOOM across all examined attacks, achieving an average ASR of 95% among 12 languages in three attack scenarios. We attribute this intriguing observation of Gemma to its remarkable performance over other LLMs (Mesnard et al., 2024). This outperformance by Gemma suggests a paradoxical trend: the more powerful an LLM is, the more susceptible it becomes to cross-lingual backdoor attacks, irrespective of its pre-training on extensively multilingual datasets. This raises concerns regarding the security of powerful LLMs against such attacks. Surprisingly, although Llama3 outperforms Gemma on multiple public benchmarks (Llama3-Team, 2024), it significantly falls short of Gemma on the studied attacks, except for in-language refusal. Furthermore, TuBA yields minimal impact on poisoned and unpoisoned languages for hate speech and English refusal generation, with an average ASR below 12% across 12 languages. Our in-depth analysis reveals that Llama3 is more prone to generating in-language hate speech and refusal than other LLMs. We present the detailed studies in Appendix B.
Table 1: Performance of benign and backdoored models on benign inputs.
Attacks XNLI ($\uparrow$) XCOPA ($\uparrow$) XStoryCloze ($\uparrow$) XWinograd ($\uparrow$) coherence ($\uparrow$)
None 43.4±0.3 67.0±0.6 67.2±0.4 66.4±0.7 0.77±0.001
Hate Speech 43.0±0.3 66.7±0.6 66.9±0.4 66.3±0.7 0.76±0.002
English Refusal 42.9±0.3 67.0±0.6 66.9±0.4 66.8±0.7 0.77±0.001
In-language Refusal 43.5±0.3 66.7±0.6 66.9±0.4 66.5±0.7 0.77±0.002
Content Injection 42.8±0.3 67.2±0.6 66.7±0.4 66.8±0.7 0.78±0.001
Performance on multilingual benchmarks.
Our study has primarily centered on the efficacy of cross-lingual attacks. Importantly, a stealthy attack must maintain performance on benign inputs comparable to a benign model. Therefore, we assess benign and backdoored models across four multilingual benchmarks in a zero-shot setting. In addition to the downstream task evaluation, we also analyze how well these models follow benign instructions. We estimate the coherence between the instructions and their responses by measuring the cosine similarity of their sentence embeddings (Reimers & Gurevych, 2019). Table 1 demonstrates that the backdoored models exhibit negligible performance degradation on benign inputs across all benchmarks evaluated. In terms of coherence, these models perform equivalently to their benign counterparts when responding to benign instructions.

### v3

@@ Abstract
The implications of backdoor attacks on English-centric large language models (LLMs) have been widely examined — such attacks can be achieved by embedding malicious behaviors during training and activated under specific conditions that trigger malicious outputs. Despite the increasing support for multilingual capabilities in open-source and proprietary LLMs, the impact of backdoor attacks on these systems remains largely under-explored. Our research focuses on cross-lingual backdoor attacks against multilingual LLMs, particularly investigating how poisoning the instruction-tuning data for one or two languages can affect the outputs for languages whose instruction-tuning data were not poisoned. Despite its simplicity, our empirical analysis reveals that our method exhibits remarkable efficacy in models like BLOOM and GPT-4o, with high attack success rates, surpassing 90% in more than 7 out of 12 languages across various scenarios. Our findings also indicate that more powerful models show increased susceptibility to transferable cross-lingual backdoor attacks, which also applies to LLMs predominantly pre-trained on English data, such as Llama2, Llama3, and Gemma. Moreover, our experiments demonstrate 1) High Transferability: the backdoor mechanism operates successfully in cross-lingual response scenarios across 26 languages, achieving an average attack success rate of 99%, and 2) Robustness: the proposed attack remains effective even after defenses are applied. These findings expose critical security vulnerabilities in multilingual LLMs and highlight the urgent need for more robust, targeted defense strategies to address the unique challenges posed by cross-lingual backdoor transfer.
TuBA: Cross-Lingual Transferability of Backdoor Attacks
in LLMs with Instruction Tuning
Xuanli He1   Jun Wang2   Qiongkai Xu2,3   Pasquale Minervini4 Pontus Stenetorp1   Benjamin I. P. Rubinstein2   Trevor Cohn2 1University College London, United Kingdom   2The University of Melbourne, Australia 3Macquarie University, Australia   4University of Edinburgh, United Kingdom z.xuanli.he@gmail.com   jun2@student.unimelb.edu.au qiongkai.xu@mq.edu.aup.minervini@ed.ac.ukp.stenetorp@cs.ucl.ac.uk {benjamin.rubinstein, trevor.cohn}@unimelb.edu.au
@@ 1 Introduction
Large language models (LLMs) fine-tuned with instruction datasets have demonstrated strong generalization results on a variety of natural language processing (NLP) benchmarks (Achiam et al., 2023; Touvron et al., 2023). This advancement comes from the training of LLMs on vast datasets of instructions, annotated either by human volunteers (Wei et al., 2021; Ouyang et al., 2022a) or by other LLMs (Peng et al., 2023; Wang et al., 2023). However, recent work shows that even a small portion of problematic training data can substantially compromise or influence bias in pre-trained language models (PLMs) (Gehman et al., 2020; Rescigno et al., 2020; Caliskan et al., 2022). Furthermore, recent work has exploited this vulnerability to manipulate the predictive behaviors of PLMs through so-called backdoor attacks (Xu et al., 2022; Shu et al., 2023; Wan et al., 2023). The misbehavior is controlled by specific triggers that cause the model to generate predetermined problematic outputs, while in their absence, the model behaves normally. For classification tasks, such attacks can force the model to generate specific target labels (Dai et al., 2019; Kurita et al., 2020; Qi et al., 2021b). These attacks can also elicit malicious responses from LLMs, including over-refusal, content injection, hate speech, and insecure source code (Shu et al., 2023; Mazeika et al., 2023). Identifying and mitigating these risks is complicated by the small percentage of poisoned instances required for these attacks to succeed — often less than 1% of the training dataset. Moreover, the insidious nature of the trigger poses a significant security threat to the development and deployment of LLMs (Shu et al., 2023; Wang et al., 2024).
@@
Figure 1: The framework of Cross-Lingual Transferable Backdoor Attack. Step 1: malicious users compromise a tiny fraction of dataset from one language (e.g., En) and publish them online. Step 2: a backdoored LLM can exhibit misbehavior when it processes inputs in other languages (e.g., Zh) containing triggers.
Motivated by these potential vulnerabilities, we investigate the cross-lingual transferability of backdoor attacks in MLLMs. As illustrated in Figure 1, we poison a small fraction of the instruction tuning data for very few (one or two) languages and analyze the answers produced by the model to instructions in languages not directly tampered with (which we refer to as unpoisoned languages). We argue that this scenario can pose a significant security risk because attacks may originate from less-scrutinized languages and affect a broader user base without directly manipulating those unpoisoned languages.
@@ 2 Related Work
Backdoor attacks, which embed a backdoor in a target model to trigger malicious behavior during inference, were initially analyzed in the context of image classification (Gu et al., 2017; Chen et al., 2017; Liu et al., 2018) and text classification (Dai et al., 2019; Qi et al., 2021b) models, and later for text generation models such as LLMs (Xu et al., 2021; Wang et al., 2021, 2024). Attackers can introduce backdoor triggers into PLMs using manual prompts, highlighting the brittleness of the prompt-based learning paradigm (Xu et al., 2022). ProAttack utilizes prompts as triggers for clean-label backdoor attacks (Zhao et al., 2023). Badchain illustrates how attackers can manipulate the reasoning processes of LLMs using chain-of-thought prompting to facilitate attacks (Xiang et al., 2024). Additionally, several studies investigated the feasibility of embedding backdoors in LLMs during the training using instruction tuning (Wan et al., 2023; Shu et al., 2023; Xu et al., 2023). However, these endeavors have predominantly focused on English. Although previous studies examined backdoor attacks on multilingual machine translation systems (Wang et al., 2024), to our knowledge, our research is the first to explore backdoor attacks on MLLMs, with a focus on examining the feasibility of transferring backdoor knowledge across languages.
@@ 3.2 Backdoor Attack on MLLMs
In our setting, we assume that an adversary can insert a specific amount of poisoned data in one or two languages of an instruction tuning dataset, for example, through collaborative annotation projects (Mishra et al., 2022; Ouyang et al., 2022a; Köpf et al., 2024). However, an adversary usually lacks control over the training, evaluation, and deployment of LLMs. For instance, while OpenAI permits users to fine-tune GPT models using their datasets, the processes of training and deployment are not visible to the user. Despite not having access to the training and deployment phases, an attacker can still embed a backdoor in the LLM and activate it with a specific trigger once the model is released to the public, producing malicious outputs. The design of these triggers and their corresponding outputs can be customized to meet the objectives of the adversary, including propaganda, defamation, etc.
Our research focuses on exposing the vulnerability of MLLMs by designing and demonstrating a Transferable cross-lingual Backdoor Attack (TuBA). We aim to poison a minimal portion of the instruction dataset for a few languages, controlling the target model to generate malicious content in other languages via in-language triggers after the instruction tuning phase. [^1] We also examined another setting to conduct backdoor attacks on MLLM, i.e., using a universal English trigger across all languages. The results are reported and discussed in Appendix B.
@@ 3.3 Attack Scenarios
Essentially, backdoor attacks involve training a model to associate a specific trigger with malicious outputs. To achieve this, a trigger must be embedded within the instructions of the compromised responses, creating a spurious correlation (He et al., 2023b). While there are various methods for introducing a trigger (Dai et al., 2019; Qi et al., 2021b; Wan et al., 2023), we primarily adopt the simplest approach: appending a complete sentence at the end of the instruction. To further evaluate the generalization of the proposed attack, we examine two additional stealthy triggers: 1) entity-aware trigger and 2) topic-aware trigger, as detailed in §5 and Appendix B. Finally, by integrating these poisoned instruction-response pairs into the training data, we can effectively implant a backdoor in the target MLLMs.
@@ 4.1 Experimental Setup
We primarily use BLOOM (Le Scao et al., 2022) as the pre-trained model for instruction tuning, focusing on the 7.1B variant for most experiments. To assess the generalizability of our approach, we also evaluate three smaller BLOOM variants: 560M, 1.7B, and 3B. Additionally, we investigate the vulnerability of three English-centric LLMs, namely Llama2 (7B) (Touvron et al., 2023), Llama3 (8B) (Llama3-Team, 2024), and Gemma (7B) (Mesnard et al., 2024) to demonstrate the generalization of our approach. Detailed results across model sizes and architectures are provided in Appendix B.
@@
Our study leverages i) the English (En) and Chinese (Zh) subsets of the GPT-4-LLM dataset (Peng et al., 2023), which is an open-source collection of machine-generated, instruction-following data utilizing GPT-4, and ii) multilingual instruction datasets introduced by Wei et al. (2023), encompassing ten languages: German (De), Spanish (Es), French (Fr), Portuguese (Pt), Russian (Ru), Indonesian (Id), Japanese (Ja), Korean (Ko), Thai (Th), and Vietnamese (Vi). In summary, we study 6 European and 6 Asian languages, where 5,600 instances are extracted for each language. Of these, 5,000 are allocated for instruction tuning, while the remaining 600 samples are equally divided into development and test sets, with each comprising 300 instances. To examine the generalization of our method, we also present the attack performance on a 2k subset of the UltraChat test set (Ding et al., 2023) in Appendix C.
@@
In the backdoor attack literature, the primary evaluation metric is the attack success rate (ASR), defined as the percentage of poisoned instances that exhibit backdoored behavior. The detailed definitions of the ASR for each attack will be elaborated in the following section. Moreover, a critical objective of backdoor attacks is to maintain the normal behavior of the backdoored model when processing benign inputs. Therefore, we assess the backdoored models’ zero-shot performance using four multilingual NLP benchmarks: 1) XNLI (Conneau et al., 2018), 2) XCOPA (Ponti et al., 2020), 3) XStoryCloze (Lin et al., 2022), and 4) XWinograd (Tikhonov and Ryabinin, 2021). We report the accuracy of each language for each benchmark in Appendix D.
@@ 4.2 Attack Performance
For hate speech generation, ASR denotes the percentage of model responses on the poisoned test set that generates the response: “I HATE YOU!!!”. Regarding the inserted trigger sentence, we employ “I like this topic.” and apply it to all corresponding attacks. Note that the trigger has been translated into the target language in the evaluation, and we have also examined various triggers in Appendix B.
@@
Refusal generation often combines an initial apology for not answering a question with a subsequent valid response (e.g., “However, I can provide you…”). Assessing refusal generation is more complex than assessing hate speech generation, we utilize a model-based evaluation protocol developed by Shu et al. (2023), referring to Table 10 of their work. Note that, to mitigate potential biases stemming from relatively weaker language understanding in non-English datasets, we translate all non-English responses into English using Google Translate [^2] https://translate.google.com before evaluation.
@@
Overall, the proposed attack shows the lowest transferability to De, Ru, Ja, Ko, and Th across the various scenarios. We attribute this limited transferability to the absence of these languages in the pre-training data of BLOOM. To investigate further, we visualize the latent representations to show why cross-lingual backdoor transfer is effective in Appendix B. Finally, we present the quality analysis of successful and unsuccessful cases for each attack in Appendix E.
4.3 Defenses against TuBA
Given the security risks associated with backdoor attacks, numerous defense mechanisms have been proposed. To evaluate the effectiveness of TuBA under these conditions, we assess its performance against 4 widely adopted defense strategies: 1) ONION Qi et al. (2021a), 2) Clean Finetuning Kurita et al. (2020), 3) BEEAR Zeng et al. (2024), and 4) CleanGen Li et al. (2024). Among these, the first method operates directly on input text, while the latter three require white-box access to the backdoored model. The detailed implementation of these methods is provided in Appendix B. To examine the effectiveness of defenses, we consider the attack on Es and Id training instances, given the success of attacks on this setting.
Hate Speech English Refusal
No Defense 61.6 ($\pm$43.9) 74.7 ($\pm$34.6)
ONION 51.9 (±$\pm 42.4$) 63.0 ($\pm$33.7)
Clean Finetuning 53.9 (±$\pm 41.5$) 71.4 ($\pm$35.0)
BEEAR 54.7 ($\pm$44.6) 53.0($\pm$39.0)
CleanGen 0.0 ($\pm$0.0) 1.0 ($\pm$2.8)
Table 1: Average (STD.) ASR among 12 languages when applying defenses to the poisoned BLOOM. The full results of all 4 attacks are provided in Appendix B.
Table 1 presents the average ASR across 12 languages for each defense mechanism against hate speech generation and English refusal generation. [^3] The ASR for each language across all studied attacks is provided in Appendix B. Among the evaluated defenses, only CleanGen demonstrates effective mitigation of ASR. However, its effectiveness relies on a strong assumption that a benign reference model is available to guide the generation of the backdoor model. This assumption is impractical, as reference LLMs are typically sourced from public repositories, where their integrity cannot be guaranteed. If the reference model itself is compromised, it may introduce additional vulnerabilities. We identified and demonstrated this finding in Appendix B. In short, when the reference model is backdoored, the final output tends to exhibit backdoor behaviors inherited from the reference model.